OpenStudy (anonymous):

Juan recorded his brother bungee jumping from a height of 1100 feet. At the time the cord lifted his brother back up, he was 76 feet above the ground. If Juan started recording as soon as his brother fell, how much time elapsed when the cord snapped back? Use f(t)=-16t^2+c, where c is the height.

OpenStudy (anonymous):

f(t)= -16t^2+1100 at t=x f(x) = 76 so solve for x 76=-16x^2+1100

OpenStudy (anonymous):

I got x = 8i?

OpenStudy (anonymous):

double check your math

OpenStudy (anonymous):

76=-16x^2+1100 -1100+76=-16x^2 -1024=-16x^2 divide by -16 64=x^2

OpenStudy (anonymous):

I did: 76=-16x^2+1100 -1100 -1100 -1024=16x^2 /16 /16 -64 = x^2 -8i = x

OpenStudy (anonymous):

you cant just get rid of the negative sign in front of the 16

OpenStudy (anonymous):

Oh crap i totally didnt see the - sign. My bad. So the answer would just be x = 8?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

no such thing as imaginary time

OpenStudy (anonymous):

Thank you so much! and for explaining it!

OpenStudy (anonymous):

no prob